# Efficient Nearest Neighbor Search Using Dynamic Programming

### v3

@@ VI Desirable Properties
Beyond its speed advantages, we believe the current method holds significant potential, and future high-quality research could further expand the algorithm’s capabilities. In the following section, we will analyze several key aspects.
VI-A Nearest Neighbor Among the First $n$ Points
@@
$\rho_{i}$ is defined as the density of point $p_{i}$, and $d_{ij}$ represents the distance between points $p_{i}$ and $p_{j}$. Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating the distance between every pair of points, which has a complexity of $O(n^{2})$. However, for 3D points, using our algorithm, this problem can be solved in $O(n\log n)$ time.
@@
Farthest Point Sampling (FPS) is a widely used sampling algorithm due to its ability to ensure uniform sampling of points, making it broadly applicable. For instance, in the PointNet++ [25] framework for 3D point cloud deep learning, FPS is employed to sample and cluster points as part of the receptive field. The complexity of sampling $k$ points from $n$ points is $O(kn)$. Currently, there are approximate farthest point sampling algorithms designed for specific types of point clouds [26], but as far as we know, no breakthroughs have been made in reducing the complexity of exact farthest point sampling. However, based on our algorithm, we achieve a complexity of $O(k\log n)$ for sampling $k$ points from $n$ points in 3D space.
@@ VI-C Various Primitives & Distance Metrics
VI-D Fast Preprocessing & Slow Query
We acknowledge that the preprocessing efficiency of our algorithm is relatively low, especially in high-dimensional spaces. Furthermore, as previously mentioned, the incremental construction of Voronoi diagrams for various primitives under different distance metrics presents significant challenges. Therefore, we propose a new direction. As discussed in Section IV-A, when constructing the Query Table, we need to identify the set $\omega_{m}$ upon the insertion of $p_{m}$, as shown in Algorithm 1. The set is defined as follows:
@@ VI-D Fast Preprocessing & Slow Query
In fact, when dealing with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes extremely challenging. Therefore, we compute $\omega_{m}^{*}$ instead, which only needs to satisfy $\omega\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results. This approach offers a direction for performing fast nearest neighbor searches in high-dimensional spaces and scenarios involving different primitives and distance metrics. As a result, the computed set may contain redundancies, leading to a slight reduction in the speed of the nearest neighbor search. Additionally, we believe that even if $\omega_{m}$ does not belong to $\omega_{m}^{*}$, it could provide valuable insights for developing approximate nearest neighbor search algorithms.
VI-E Fast Query & Slow Preprocessing
@@
As illustrated in Figure 12, (a) and (b) depict a simple incremental construction process, with the red dot representing the query point $q$. In this example, we have only utilized the information shown in (c), namely that $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$. As shown in brown in (c). However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$. As shwon in (d). This enables us to design a more rigorous and precise occupancy evaluation. Our method currently does not leverage this information during the query process, even though it can be obtained through complex preprocessing, which would, of course, involve substantial preprocessing overhead.
@@
Figure 12: (a) and (d): A simple incremental insertion process, with the red dot representing the query point $q$. (c) Based on the query process of (a)(b), we know the information $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown. (d) In fact, the complete information we can obtain is $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
However, we also recognize that utilizing all of the aforementioned information would result in significantly increased preprocessing complexity. To balance this, we implemented a simplified approach that compromises between the algorithm described in this paper and the information mentioned in this section. In brief, in addition to $q\in\text{Cell}(p_{i};\mathcal{V}_{i})$, we incorporated $q\in\text{Cell}(\Phi_{i-1};\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(p_{i};\mathcal{V}_{i})$. This region is used as the fundamental unit for occupancy evaluation, rather than a complete Voronoi cell.
We were pleasantly surprised to find that, despite implementing only a simplified version, we achieved a significant reduction in the number of comparisons compared to Table I. As shown in Table IV, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section. Remarkably, only around 100 pairwise distance comparisons are required to find the exact nearest neighbor for a query point among 1 million points. However, we must acknowledge that while the number of comparisons has decreased, the actual query time has increased. We speculate that this may be due to increased memory usage by the data structure during the query phase, which likely results in more cache misses, thereby reducing efficiency. Nevertheless, we believe this is a highly meaningful direction for further exploration.
TABLE IV: The number of pairwise comparisons during the query process using different construction strategies.
@@ VII Limitations and Future Work
The main drawback of our algorithm at present is its preprocessing cost, which limits its application in some scenarios. However, fortunately, our method supports nearest neighbor queries during the construction process, which may assist in the tetrahedron location problem during incremental Delaunay construction, offering a potential avenue for accelerating preprocessing. Therefore, in future work, we will focus on improving the preprocessing speed. Additionally, our method currently only supports nearest neighbor search; we aim to extend the current algorithm to support k-nearest neighbor queries in the future. Furthermore, all the properties mentioned in Section VI are goals we aim to address in our future work.
@@ VIII Conclusion
In this paper, we observe that when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points, the efficiency of traditional nearest neighbor search algorithms decreases. In certain special cases, the complexity can degrade to $O(n)$. Inspired by the incremental Delaunay construction process, we propose a nearest neighbor search algorithm specifically for 2D maniflod point cloud. Our method achieves 1-10 times the speed of KD Tree in nearest neighbor searches. We validated the efficiency advantages of our approach through extensive experiments.
Additionally, our algorithm demonstrates significant potential. For instance, it supports querying the closest point to $q$ among any of the first $k$ points. This capability facilitates the application of peak clustering algorithms to large-scale data scenarios. Moreover, we have implemented a farthest point sampling algorithm with $O(k\log n)$ complexity in 3D space. Our mathod can efficiently identify the point closest to the query point from a set of 1 million points by performing about 100 pairwise distance comparisons.
We believe that the algorithm presented in this paper holds significant potential for application and will make a substantial contribution to the community.
[…]

### v4

@@ VI Desirable Properties
In light of its speed advantages, we believe that the current method harbors substantial potential for further advancements. In the following, we will delineate several pivotal aspects pertaining to its applications, aiming to elucidate the breadth and depth of its prospective utility in various domains.
VI-A Nearest Neighbor Among the First $k$ Points
@@
$\rho_{i}$ is defined as the density of point $p_{i}$, and $d_{ij}$ represents the distance between points $p_{i}$ and $p_{j}$. Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating the distance between every pair of points, which has a complexity of $O(n^{2})$. In contrast, our algorithm can efficiently solve this problem in $O(n\log n)$ time, assuming the data points are two-dimensional or three-dimensional. However, for higher-dimensional data, the time complexity of our algorithm increases due to the computationally expensive Delaunay triangulation process.
@@
Farthest Point Sampling (FPS) is a widely used sampling algorithm because it ensures uniform point distribution, making it highly applicable across various tasks. For instance, FPS is used in the PointNet++ [35] framework for 3D point cloud deep learning to ensure uniform sampling and clustering of points, which contributes to building an effective receptive field. Although the initial purpose of our method was not specifically for farthest point sampling, it naturally lends itself to a simple and fast FPS strategy. Compared to existing methods, our proposed FPS strategy is more straightforward and easier to implement, offering an intuitive solution for efficient point sampling.
@@ VI-C Various Primitives & Distance Metrics
VI-D Fast Preprocessing with Compromising Query Performance
In what follows, we aim to enhance the preprocessing efficiency at the expense of query performance. As elaborated in Section IV-A, during the construction of the Query Table, it is essential to identify the set $\omega_{m}$ upon inserting $p_{m}$, as illustrated in Algorithm 1. The set is defined by:
@@ VI-D Fast Preprocessing with Compromising Query Performance
Indeed, when confronted with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes exceedingly challenging. Consequently, we opt to compute $\omega_{m}^{*}$, which merely needs to satisfy $\omega_{m}\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results. This approach opens up avenues for executing rapid nearest neighbor searches in high-dimensional spaces and diverse primitive and distance metric contexts. As an inevitable outcome, the computed set may harbor redundancies, thereby causing a marginal decrease in the speed of the nearest neighbor search.
Moreover, even if $\omega_{m}$ does not fall within $\omega_{m}^{*}$, it could furnish valuable insights conducive to the development of approximate nearest neighbor search algorithms.
VI-E Fast Query with Elaborated Preprocessing
@@
As illustrated in Figure 14, (a) and (b) depict a simple incremental construction process, with the red dot representing the query point $q$. In this example, we have only utilized the information shown in (c), namely that $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$. As shown in brown in (c). However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as is shown in (d). This enables us to design a more rigorous and precise evaluation criterion. Our method currently does not leverage this information during the query process, even though it can be obtained through complex preprocessing, which would, of course, involve substantial preprocessing overhead.
@@
Figure 14: (a) and (b): A simple incremental insertion process, with the red dot representing the query point $q$. (c) Based on the query process of (a)(b), we know the information $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown. (d) In fact, the complete information we can obtain is $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
However, we also recognize that utilizing all of the aforementioned information would result in significantly increased preprocessing complexity. To balance this, we implemented a simplified approach that compromises between the algorithm described in this paper and the information mentioned in this section. In brief, in addition to $q\in\text{Cell}(p_{i};\mathcal{V}_{i})$, we incorporated $q\in\text{Cell}(\Phi_{i-1};\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(p_{i};\mathcal{V}_{i})$. This region is used as the fundamental unit for occupancy evaluation, rather than a whole Voronoi cell.
We were pleasantly surprised to find that, despite implementing only a simplified version, we achieved a significant reduction in the number of comparisons compared to Table I. As shown in Table V, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section. Remarkably, only around 100 pairwise distance comparisons are required to find the exact nearest neighbor for a query point among 1 million points. However, we must acknowledge that while the number of comparisons has decreased, the actual query time has increased. We speculate that this may be due to increased memory usage by the data structure during the query phase, which likely results in more cache misses, thereby reducing efficiency. Nevertheless, we believe this is a highly meaningful direction for further exploration.
TABLE V: The number of pairwise comparisons during the query process using different construction strategies.
@@ VII Limitations and Future Work
The primary limitation of our algorithm is its high preprocessing cost, which restricts its applicability, particularly in large-scale scenarios or with high-dimensional datasets where efficiency is crucial. Therefore, in future work, we plan to couple the construction of our algorithm more closely with Delaunay computations, or alternatively, develop an approximate Delaunay construction to achieve faster preprocessing. These approaches aim to reduce the preprocessing burden and improve scalability for large-scale and high-dimensional applications.
Additionally, our method currently only supports nearest neighbor search. In future work, we aim to extend the algorithm to support k-nearest neighbor queries, which would enhance its versatility. Moreover, addressing all the desirable properties outlined in Section VI remains a significant long-term goal for us.
@@ VIII Conclusion
In this paper, we observe that traditional nearest neighbor search algorithms exhibit decreased efficiency when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points. In certain special cases, the complexity can degrade to $O(n)$. Inspired by the incremental Delaunay construction process, we propose a novel nearest neighbor search algorithm specifically designed for 2D manifold point cloud data. Our method achieves a speedup of 1-10 times compared to the KD Tree approach. We validated the efficiency advantages of our approach through extensive experiments. The high performance of our algorithm is also supported by evidence showing that it requires only about 100 pairwise distance comparisons to identify the point closest to the query point from a set of 1 million points.
Additionally, our algorithm demonstrates significant potential. For instance, it supports querying the closest point to $q$ among any of the first $k$ points. This capability facilitates the application of peak clustering algorithms to large-scale data scenarios. Moreover, we have implemented a farthest point sampling algorithm with $O(k\log n)$ complexity in 3D space. We believe that the algorithm presented in this paper holds significant potential for application and will make a substantial contribution to the community.